OpenStudy (amenah8):

Find the indeterminate form and locate the limit using l'hopital's rule: Find the limit as x approaches INFINITY at (ln(2x)-ln(x+1))

OpenStudy (freckles):

use log properties

OpenStudy (freckles):

log(a)-log(b)=log(a/b)

OpenStudy (amenah8):

log(2x)-log(x+1)?

OpenStudy (freckles):

use the law I just mentioned

OpenStudy (freckles):

it works for any base

OpenStudy (freckles):

ln(2x)-ln(x+1)=?

OpenStudy (freckles):

also notifications aren't working too well for me

OpenStudy (amenah8):

so log(2x)-log(x+1)=log2x/(x+1)

OpenStudy (freckles):

\[\ln(2x)-\ln(x+1)=\ln(\frac{2x}{x+1})\]

OpenStudy (freckles):

take the limit inside the the log there

OpenStudy (amenah8):

what do you mean by take the limit inside the log?

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty} \ln(\frac{2x}{x+1})=\ln(\lim_{x \rightarrow \infty} \frac{2x}{x+1})\]

OpenStudy (amenah8):

so now i put in infinity for x?

OpenStudy (freckles):

yes but you get inf/inf and the directions wants you to find the limit using l'hospital and you can since you have inf/inf

OpenStudy (amenah8):

so now take the derivative of the top and bottom

OpenStudy (amenah8):

2

OpenStudy (freckles):

yes so you have ln(2) as the limit